# A First Look at Bugs in LLM Inference Engines

### v1

@@ Abstract.
Large language model-specific inference engines (in short as LLM inference engines) have become a fundamental component of modern AI infrastructure, enabling the deployment of LLM-powered applications (LLM apps) across cloud and local devices. Despite their critical role, LLM inference engines are prone to bugs due to the immense resource demands of LLMs and the complexities of cross-platform compatibility. However, a systematic understanding of these bugs remains lacking. To bridge this gap, we present the first empirical study on bugs in LLM inference engines. We mine official repositories of 5 widely adopted LLM inference engines, constructing a comprehensive dataset of 929 real-world bugs. Through a rigorous open coding process, we analyze these bugs to uncover their symptoms, root causes, and commonality. Our findings reveal six major bug symptoms and a taxonomy of 28 root causes, shedding light on the key challenges in bug detection and location within LLM inference engines. Based on these insights, we propose a series of actionable implications for researchers, inference engine vendors, and LLM app developers.
copyright: nonejournalyear: 2025
@@ 1. Introduction
Large language model-specific inference engine (abbreviated as LLM inference engine) has emerged as a vital component of modern artificial intelligence infrastructure. These engines are specially optimized for LLM inference, offering LLM-specific optimizations tailored to reduce resource consumption and enhance inference speed. LLM inference engines have significantly boosted the adoption of LLM-powered applications (LLM apps). While LLM inference engines deliver fast, high-quality responses in cloud environments (cha, 2024), they also support local deployment on personal devices like Windows, Linux, and MacOS PCs for personal LLM apps such as Ollama (oll, 2024).
LLM inference engine is complexly constructed and highly bug-prone due to the immense resource demands of LLMs and the challenges of cross-platform compatibility. First, due to the vast number of parameters in LLMs, LLM inference is both compute-intensive and memory-hungry. To enhance inference efficiency, LLM inference engines employ numerous complex optimizations to reduce the memory consumption and enhance the inference latency, such as paged attention, continuous batching, and speculative decoding. Second, the engines should adapt to a wide range of application settings. They need to support various LLM architectures across different environments, including multi-node clusters, single nodes with multi-GPUs, single GPU servers, PCs, mobile devices, IoT devices, and browsers. The combination of the complex optimizations and diverse environment adoption potentially leads to more bugs, which can severely impact the availability, performance, and user experience of LLM apps.
Bugs in AI frameworks have been extensively studied. Previous work has empirically investigated bugs of deep learning (DL) frameworks across stages, including the learning stage (Islam et al., 2019; Zhang et al., 2018, 2020; Humbatova et al., 2020), the compiling stage (Shen et al., 2021; Guan et al., 2023), and the deployment stage (Chen et al., 2020, 2021). However, a systematic understanding and characterization of bugs in LLM inference engines remains lacking. This knowledge is urgently needed as LLMs have become the dominant AI paradigm.
To bridge this gap, we present the first empirical study on bugs in LLM inference engines. Given the increasing prevalence of LLM apps, this research is of great significance. Specifically, we aim to address three research questions (RQs):
@@
To answer the RQs, we carefully select five representative LLM inference engines for our analysis, including Llama.cpp (lla, 2024b), vLLM (vll, 2025), DeepSpeed (dee, 2025), Mlc-llm (team, 2023), and TensorRT-llm (ten, 2025). These engines are commonly used in real-world LLM apps and exhibit diversity in their implementations, ensuring the validity and generalizability of our results. We gather all the bug issues from the official repositories of these engines as of December 1, 2024. Following the open coding procedure (Seaman, 1999), we review error reports, source code, developer discussions, and related pull requests (PRs) of issues to uncover the symptoms, root causes, and distribution of bugs, constructing a dataset with 929 manually labeled real-world bugs. Based on the dataset, we identify six symptom types and construct a taxonomy of 28 specific root causes. Our main findings include that (1) Under undetermined LLM output, five factors (i.e., mis-encoded characters, incoherent semantics, inconsistent outputs, abnormal lengths, and repeated substrings) provide insights for oracle setting; (2) Testing LLM inference engines requires diverse oracles beyond mere crash detection, as over 35% of bugs manifest as non-crash symptoms. (3) We identify actionable diagnostic patterns for the observed bug symptoms. For instance, crashes, feature failures, and system hangs necessitate checks on functionality, environment, and configuration. (4) Bug symptoms and root causes show a strong correlation across engines, with correlation coefficient greater than 0.5, indicating common bug patterns despite different engine implementations.
@@ 2.1. LLM Deployment
LLM apps have gained widespread use, such as cloud-based chatbots (cha, 2024), and personal AI assistants deployed on user devices (app, 2024; goo, 2024). In these apps, LLM inference engines are crucial components for enhancing inference efficiency and adapting LLMs to various environments. Generally, the workflow of LLM deployment falls into three phases, including engine setup, model conversion, and inference/serving, as shown in Fig. 1.
@@ 2.2. LLM Inference Engine
Figure 2. Architecture of LLM inference engine
The primary design goals of LLM inference engines are compatibility and efficiency. First, compatibility ensures that various series of LLM run smoothly across diverse environments. Engines like vLLM (vll, 2025) and DeepSpeed (dee, 2025) are tailored for diverse cloud-based distributed inference scenarios, while engines like Llama.cpp (lla, 2024b) and Mlc-llm (team, 2023) focus on compatibility with heterogeneous personal devices and backends. Second, efficiency optimization focuses on integrating algorithms that maximize resource utilization and minimize inference latency under varying resource conditions. Engines generally achieve efficiency by optimizing model conversion, inference strategies, operator implementation, and resource management.
As shown in Fig. 2, LLM inference engines typically support dozens of popular LLM series, such as Llama (lla, 2024a) and Gemma (gem, 2024). Some of the engines (lla, 2024b; team, 2023; ten, 2025) provide generally supported model formats and conversion tools for user-customized LLMs. They accommodate a wide range of devices, including distributed servers, single-GPU servers, PCs, smartphones, and IoT devices. Some of them (lla, 2024b; team, 2023) are also compatible with various platforms like Linux, Windows, and Android, as well as backends such as CUDA, Vulkan, and Metal. Next, we discuss the major components of the engines.
These engines often provide setup tools for custom engine construction and deployment in the engine setup phase and offer model conversion tools for the model conversion phase. In the inference/serving phase, users can interact with LLMs via command lines or inference services. Upon receiving a start command, the engine initializes the LLM inference service based on the configurations and loads the necessary LLMs onto the selected backend. When the service receives a user request with specific prompts and job configurations, the engine allocates resources, creates a job, checks the chosen model, and performs auto-regressive token generation to respond to users. This process involves several key components, as shown in Fig. 2.
The engine initializer processes service configurations to set up the inference environment, including loading dependent modules, configuring parameters, coordinating with the model loader to prepare LLMs, and opening the service port to await inference requests. The model loader is responsible for loading the specified converted LLM onto the chosen backend.
@@ 3.1. Engine Selection
We begin with a widely recognized leaderboard of state-of-the-art LLM training and inference projects actively maintained by the community (awe, 2025). Since our study focuses on inference-related bugs, we first exclude engines dedicated to LLM training. We also remove closed-source engines due to their lack of accessible bug reports. From the remaining candidates, we select the top five LLM inference engines based on the number of stars in their official GitHub repositories, a widely used indicator of popularity (Liu et al., 2024; Li et al., 2020). As a result, we adopt Llama.cpp (lla, 2024b), vLLM (vll, 2025), DeepSpeed (dee, 2025), MLC-LLM (team, 2023), and TensorRT-LLM (ten, 2025). The details of these engines are summarized in Table 1.
Table 1. Details of selected LLM inference engines.
Engine Provider Targeted Env. GitHub Stars
Llama.cpp Ggml Universal 76,157
vLLM UCB Cloud 40,876
DeepSpeed Microsoft Cloud 37,282
Mlc-llm CMU Universal 20,150
TensorRT-llm Nvidia Cloud/Edge 9,641
These engines are representative for several reasons. First, they have received between 9,641 and 76,157 GitHub stars, indicating strong popularity. Second, they have been adopted by major organizations. For example, vLLM serves as the backend for Hugging Face, DeepSpeed is integrated into Microsoft’s infrastructure, and TensorRT-LLM is utilized by Nvidia. Third, they show diversity in targeted environments. Llama.cpp and MLC-LLM are designed for universal deployment, supporting various operating systems (such as Linux, Windows, macOS) and platforms, including PCs, mobile devices, and IoT devices. vLLM and DeepSpeed are optimized for cloud environments, while TensorRT-LLM supports both cloud-based deployments and edge servers. Finally, these engines incorporate cutting-edge features in LLM inference, such as continuous batching, paged attention, speculative decoding, and machine learning compilation.
@@ 3.2. Bug Collection
To collect bugs from the selected LLM inference engines, we focus on their official GitHub repositories. Specifically, we first follow previous work (Chen et al., 2023; Zhang et al., 2018) to leverage the GitHub issue tracking system to gather potential bug reports for each engine. We then refine the collected data and exclude invalid reports.
@@ 3.2.2. Refining Dataset
We observe a significant number of false positives in the collected bug reports, as the bug-related tags can be misapplied. To address this, manual filtering is necessary to remove these false positives. However, manually analyzing all the issues is a challenging task. Therefore, we adopt a sampling approach following previous work (Zhang et al., 2019; Chen et al., 2020), randomly selecting a subset of the bug reports while ensuring a 95% confidence level and a 5% confidence interval. This sampling process is applied separately to each engine, resulting in 125, 292, 126, 227, and 188 bug reports for Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively.
After sampling, we manually refine the dataset by excluding reports that meet any of the following criteria:
@@
After this manual refinement, we are left with a dataset of 929 real-world bugs related to LLM inference engines. The final counts are 124, 290, 121, 214, and 180 reports for Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively.
@@ 3.3. Manual Labeling
We manually analyze the 929 bugs to identify their symptoms and root causes, constructing corresponding taxonomies for LLM inference engine bugs. Following previous work (Chen et al., 2021; Guan et al., 2023), the manual analysis involves two key steps: initially constructing the taxonomies through a pilot labeling process, followed by a reliability analysis to evaluate the taxonomies. The entire manual analysis takes over 500 man-hours.
@@ 3.3.1. Pilot Labeling
First, two authors, each with over two years of experience in LLM-based application development, act as inspectors and jointly develop pilot taxonomies for bug symptoms and root causes. To fully understand the content of each issue, they review the title, description, code snippets, comments, answers, labels, associated PRs, and any referenced URLs. They then assign brief phrases as initial codes to capture the key features of each issue.
For symptoms, the inspectors focus on what the bug reporter observes and at which stage of LLM deployment the bug occurs. This information is often found in the issue descriptions and call stacks provided by reporters. For example, a reporter states that ‘loading a gptbigcode model that has been quantized with gptq fails’ in issue #8116 of vLLM. The inspectors classify the symptom as a crash occurring in the inference/serving stage.
In contrast, identifying root causes is more challenging and often requires in-depth analysis. Inspectors determine which part of the code failed and how by examining call stacks, bug reproduction experiments, confirmed comments from developers, issue descriptions, and associated PRs. If the relevant code is not within a modular component, they infer its component based on its functionality. For instance, in issue #3821 of vLLM, the associated PR #4021 states ‘avoid allocating a CUDA context for every GPU’, and modifies the allocation of the cache dictionary in its commits. Based on this, inspectors conclude that the bug originated from incorrect allocation in the engine’s resource manager.
After generating the initial codes for the issues, inspectors group similar codes into categories. The grouping process is iterative, with inspectors continuously reviewing the categories and issues to refine the taxonomies. Bugs associated with multiple symptoms or root causes are assigned to all relevant categories. In case of disagreements, another author, who has over two years of LLM inference engine deployment experience, acts as an arbitrator and joins the discussion to help reach consensus. The arbitrator, as well as two inspectors, finally approve all categories in the symptom taxonomy and the root cause taxonomy.

### v2

@@ Abstract.
Large language model-specific inference engines (in short as LLM inference engines) have become a fundamental component of modern AI infrastructure, enabling the deployment of LLM-powered applications (LLM apps) across cloud and local devices. Despite their critical role, LLM inference engines are prone to bugs due to the immense resource demands of LLMs and the complexities of cross-platform compatibility. However, a systematic understanding of these bugs remains lacking. To bridge this gap, we present the first empirical study on bugs in LLM inference engines. We mine official repositories of 5 widely adopted LLM inference engines, constructing a comprehensive dataset of 929 real-world bugs. Through a rigorous open coding process, we analyze these bugs to uncover their symptoms, root causes, commonality, fix effort, fix strategies, and temporal evolution. Our findings reveal six bug symptom types and a taxonomy of 28 root causes, shedding light on the key challenges in bug detection and location within LLM inference engines. Based on these insights, we propose a series of actionable implications for researchers, inference engine vendors, and LLM app developers, along with general guidelines for developing LLM inference engines.
Large Language Model, LLM Inference Engine, Bug
copyright: nonejournalyear: 2025ccs: General and reference Empirical studiesccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Large language model-specific inference engine (abbreviated as LLM inference engine) has emerged as a vital component of modern artificial intelligence infrastructure. An LLM inference engine is specialized software designed to run pre-trained LLMs, enabling their prediction or decision generation process (wik, 2019). These engines are specially optimized for LLM inference, offering LLM-specific optimizations tailored to reduce resource consumption and enhance inference speed. LLM inference engines have significantly boosted the adoption of LLM-powered applications (LLM apps). While LLM inference engines deliver fast, high-quality responses in cloud environments (cha, 2024), they also support local deployment on personal devices like Windows, Linux, and MacOS PCs for personal LLM apps such as Ollama (oll, 2024).
LLM inference engine is complexly constructed and highly bug-prone due to the immense resource demands of LLMs and the challenges of cross-platform compatibility. First, due to the vast number of parameters in LLMs, LLM inference is both compute-intensive and memory-hungry. For instance, running Llama-2-70B on a single GPU typically requires more than 80 GB of GPU memory, which makes it infeasible without model sharding or quantization techniques. Even smaller-scale engines such as Llama.cpp require more than 16 GB of RAM to execute models with 13B parameters efficiently. To enhance inference efficiency, LLM inference engines employ numerous complex optimizations to reduce the memory consumption and enhance the inference latency, such as paged attention, continuous batching, and speculative decoding. For instance, vLLM employs paged attention, which organizes the Key-Value (KV) cache into non-contiguous pages. The KV cache is a critical component for storing attention states during autoregressive generation of LLMs. This approach requires fine-grained memory management to reduce memory fragmentation caused by variable-length sequences, thereby reducing memory waste, which is a common bottleneck in LLM inference. Second, the engines should adapt to a wide range of application settings. They need to support various LLM architectures across different environments, including multi-node clusters, single nodes with multi-GPUs, single GPU servers, PCs, mobile devices, IoT devices, and browsers. The combination of the complex optimizations and diverse environment adoption potentially leads to more bugs, which can severely impact the availability, performance, and user experience of LLM apps.
Bugs in AI systems have been extensively studied. Previous work has empirically investigated bugs of deep learning (DL) systems across stages, including the learning stage (Zhang et al., 2018; Humbatova et al., 2020), the compiling stage (Shen et al., 2021; Guan et al., 2023), and the deployment stage (Chen et al., 2020, 2021). However, a systematic understanding and characterization of bugs in LLM inference engines remains lacking. This knowledge is urgently needed as LLMs have become the dominant AI paradigm and more prone to bugs compared to traditional DL systems.
To bridge this gap, we present the first empirical study on bugs in LLM inference engines. Given the increasing prevalence of LLM apps, this research is of great significance. Specifically, we aim to address the following research questions (RQs):
@@
(RQ4) Fix Effort: Which bugs require more effort to fix? Investigating fix effort quantifies the maintenance burden associated with different bugs, which could help prioritize testing and development efforts.
(RQ5) Fix Strategy: How can bugs be fixed? Identifying common bug fix strategies can guide the development of automated repair tools and inform developer guidelines.
(RQ6) Temporal Evolution: How do the distributions of bugs change over time? Analyzing temporal evolution helps identify dynamic shifts in bug patterns, offering a detailed understanding of bugs across the development progress.
To answer the RQs, we carefully select five representative LLM inference engines for our analysis, including Llama.cpp (lla, 2024b), vLLM (vll, 2025), DeepSpeed (dee, 2025), Mlc-llm (team, 2023), and TensorRT-llm (ten, 2025). These engines are commonly used in real-world LLM apps and exhibit diversity in their implementations, ensuring the validity and generalizability of our results. We gather all the bug issues from the official repositories of these engines as of December 1, 2024. Following the open coding procedure (Seaman, 1999), we review error reports, source code, developer discussions, and related pull requests (PRs) of issues to uncover the symptoms, root causes, and distribution of bugs, constructing a dataset with 929 manually labeled real-world bugs.
Based on the dataset, we identify six symptom types and construct a taxonomy of 28 specific root causes. Our main findings include that (1) Testing LLM inference engines requires diverse oracles beyond mere crash detection, as over 35% of bugs manifest as non-crash symptoms. (2) Under undetermined LLM output, five factors (i.e., mis-encoded characters, incoherent semantics, inconsistent outputs, abnormal lengths, and repeated substrings) provide insights for oracle setting to detect non-crash symptom. (3) We identify actionable diagnostic patterns for the observed bug symptoms. For instance, crashes, feature failures, and system hangs necessitate checks on functionality, environment, and configuration. (4) Bug symptoms and root causes show a strong correlation across engines, with correlation coefficient greater than 0.5, indicating common bug patterns despite different engine implementations. Based on our findings, we propose actionable implications for researchers, inference engine vendors, and LLM app developers, and offer general guidelines for developing LLM inference engines
@@ 2.1. LLM Deployment
LLM apps have gained widespread use, such as cloud-based chatbots (cha, 2024), coding assistants (Liu et al., 2024b), and personal AI assistants deployed on user devices (app, 2024; goo, 2024). In these apps, LLM inference engines are crucial components for enhancing inference efficiency and adapting LLMs to various environments. Generally, the workflow of LLM deployment falls into three phases, including engine setup, model conversion, and inference/serving, as shown in Fig. 1.
@@ 2.2. LLM Inference Engine
Figure 2. Architecture of LLM inference engine.
Design goals. The primary design goals of LLM inference engines are compatibility and efficiency. First, compatibility ensures that various series of LLM run smoothly across diverse environments. Engines like vLLM (vll, 2025) and DeepSpeed (dee, 2025) are tailored for diverse cloud-based distributed inference scenarios, while engines like Llama.cpp (lla, 2024b) and Mlc-llm (team, 2023) focus on compatibility with heterogeneous personal devices and backends. Second, efficiency optimization focuses on integrating algorithms that maximize resource utilization and minimize inference latency under varying resource conditions. Engines generally achieve efficiency by optimizing model conversion, inference strategies, operator implementation, and resource management.
Overall architecture. As shown in Fig. 2, LLM inference engines act as a middleware layer that bridges hardware environments and LLMs. At the bottom level, diverse devices (e.g., distributed servers, single-GPU servers, PCs, smartphones, and IoT devices) run on heterogeneous platforms such as Linux, Windows, and Android, and rely on different backends such as CUDA, Vulkan, and Metal. On top of these heterogeneous environments lies the LLM inference engine, which abstracts the hardware complexity and provides unified support for loading and executing models. At the upper level, these engines enable dozens of popular LLM series, including Llama (lla, 2024a) and Gemma (gem, 2024), to run seamlessly across devices and backends.
Functionality. The functionality of LLM inference engines follows the deployment workflow of three phases. In the engine setup phase, engines typically provide setup tools that allow custom construction and installation in the target environment. In the model conversion phase, they supply conversion utilities to adapt pre-trained LLMs into formats optimized for specific backends and platforms. Finally, in the inference/serving phase, users interact with LLMs through command-line interfaces or inference services. Upon receiving a start command, the engine initializes the service according to configurations, loads the converted model onto the backend, and processes user requests. For each request, the engine allocates resources, creates a job, checks the model, and performs auto-regressive token generation to produce outputs.
Key components involving inference/serving. The inference/serving stage is the most complex functionality of LLM inference engines, relying on several key components, as illustrated in Fig. 2. At the beginning of inference/serving, the engine initializer processes start configurations to set up the inference environment, including loading dependent modules, configuring parameters, coordinating with the model loader to prepare LLMs, and opening the service port to await inference requests. The model loader is responsible for loading the specified converted LLM onto the chosen backend.
@@ 3.1. Engine Selection
We begin with a widely recognized leaderboard of state-of-the-art LLM training and inference projects actively maintained by the community (awe, 2025). Since our study focuses on inference-related bugs, we first exclude engines dedicated to LLM training. We also remove closed-source engines due to their lack of accessible bug reports. From the remaining candidates, we select the top five LLM inference engines based on the number of stars in their official GitHub repositories, a widely used indicator of popularity (Liu et al., 2024a; Li et al., 2020). As a result, we adopt Llama.cpp (lla, 2024b), vLLM (vll, 2025), DeepSpeed (dee, 2025), Mlc-llm (team, 2023), and TensorRT-LLM (ten, 2025). The details of these engines are summarized in Table 1.
Table 1. Details of selected LLM inference engines (Data as of Oct. 2025). “GGML” is the underlying tensor library of Llama.cpp that enables the efficient execution of LLMs on consumer-grade hardware. “Universal” refers to the engine’s capability for broad deployment across heterogeneous operating systems and hardware.
Engine Provider Target Env. GitHub Repo Codebase Size (LoC)
Stars Commits Total C/C++ Python CUDA
Llama.cpp GGML Team Universal 87,810 6,765 230,397 194,360 18,576 17,461
vLLM UCB Cloud 60,137 10,462 458,618 7,955 413,545 37,118
DeepSpeed Microsoft Cloud 40,391 2,956 121,965 9,460 98,246 14,259
Mlc-llm CMU Universal 21,473 1,721 1,508,976 422,362 473,255 613,359
TensorRT-llm Nvidia Cloud/Edge 11,856 3,173 20,754,106 20,296,665 385,877 86,564
These engines are representative for several reasons. First, they have received over 10,000 GitHub stars, indicating strong popularity. Second, they have been adopted by major organizations. For example, vLLM serves as the backend for Hugging Face, DeepSpeed is integrated into Microsoft’s infrastructure, and TensorRT-LLM is utilized by Nvidia. Third, they show diversity in targeted environments. Llama.cpp and Mlc-llm are designed for universal deployment, supporting various operating systems (such as Linux, Windows, macOS) and platforms, including PCs, mobile devices, and IoT devices. vLLM and DeepSpeed are optimized for cloud environments, while TensorRT-LLM supports both cloud-based deployments and edge servers. Finally, these engines incorporate cutting-edge features in LLM inference, such as continuous batching, paged attention, speculative decoding, and machine learning compilation.
@@ 3.2. Bug Collection
To collect bugs from the selected LLM inference engines, we focus on their official GitHub repositories. Specifically, we first follow previous work (Chen et al., 2023; Zhang et al., 2018; Liu et al., 2023a) to leverage the GitHub issue tracking system to gather potential bug reports for each engine. We then refine the collected data and exclude invalid reports.
@@ 3.2.2. Refining Dataset
We observe false positives in the collected bug reports, as the bug-related tags can be misapplied. To address this, manual filtering is necessary to remove these false positives. However, manually analyzing all the issues is a challenging task. Therefore, we adopt a sampling approach following previous work (Zhang et al., 2019; Chen et al., 2020; Lou et al., 2020), randomly selecting a subset of the bug reports while ensuring a 95% confidence level and a 5% confidence interval. This sampling process is applied separately to each engine, resulting in 125, 292, 126, 227, and 188 bug reports for Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively.
After sampling, the first and second authors, each with over two years of experience in LLM-based application development, serve as inspectors and separately label false positives. In this process, they review the title, description, code snippets, comments, answers, and labels of each issue to determine whether the issue meets our exclusion criteria. The applied exclusion criteria include:
@@
The exclusion criteria are straightforward to identify. For instance, issues containing only the initial reports and lacking any responses can be easily identified as ’No informative comments.’ Consequently, the two inspectors encountered no conflict in identifying false positives due to the clear criteria. Furthermore, because the issues were randomly sampled, the absolute number of the excluded issues was relatively small. Specifically, they excluded 1, 2, 5, 13, and 8 false bug reports from Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively. After this manual refinement, we are left with a dataset of 929 real-world bugs related to LLM inference engines. The final counts are 124, 290, 121, 214, and 180 reports for Llama.cpp, vLLM, DeepSpeed, Mlc-llm, and TensorRT-llm, respectively.
For each issue in the refined dataset, we retrieved the creation and closure times using the GitHub API (git, 2025) to calculate the duration required for bug repair.
@@ 3.3. Manual Labeling
We manually analyze the 929 bugs to identify their symptoms and root causes, constructing corresponding taxonomies for LLM inference engine bugs. Following previous work (Chen et al., 2021; Zhang et al., 2024; Wen et al., 2021; Wang et al., 2023), the manual analysis involves two key steps: initially constructing the taxonomies through a pilot labeling process, followed by a reliability analysis to evaluate the taxonomies. The entire manual analysis takes over 500 man-hours.
@@ 3.3.1. Pilot Labeling
First, two authors, each with over two years of experience in LLM-based application development, act as inspectors and independently generate initial codes for bug symptoms and root causes. To fully understand the content of each issue, they review the title, description, code snippets, comments, answers, labels, associated PRs, and any referenced URLs. They then assign brief phrases as initial codes to capture the key features of each issue.
For symptoms, the inspectors focus on what the bug reporter observes and at which stage of LLM deployment the bug occurs. This information is often found in the issue description, call stacks, and related reproduction steps. For example, a reporter states that ‘loading a model that has been quantized fails’ in issue #8116 of vLLM. From the perspective of vLLM, the model loading is one of the steps in inferece/serving stage. Because the engine terminated unexpectedly with an unhandled exception and a stack trace during model loading (as shown in 1), preventing it from serving requests, the inspectors classify this case as a crash occurring in the inference/serving stage.
Listing 1: Error messaegs in vLLM issue #8116.
1File "xxx/gpt_bigcode.py", line 356, in load_weights
2 weight_loader(param, loaded_weight)
3File "xxx/linear.py", line 779, in weight_loader_v2
4 self._load_fused_module_from_checkpoint(param, loaded_weight)
5File "xxx/linear.py", line 762, in _load_fused_module_from_checkpoint
6 loaded_weight_shard = loaded_weight.narrow(param.output_dim,
7 ...
In contrast, identifying root causes is more challenging and often requires in-depth analysis. Inspectors determine how the code failed by examining call stacks, bug reproduction experiments, confirmed comments from developers, issue descriptions, and associated PRs. For instance, in issue #3821 of vLLM, the associated PR #4021 states ‘avoid allocating a CUDA context for every GPU’, and modifies the allocation of the cache dictionary in its commits. Based on this, inspectors infer that the bug originated from incorrect resource allocation. If a bug involves multiple root causes, inspectors assign it to all relevant root cause categories. Additionally, to investigate the vulnerable components in LLM inference engines, they also mark the associated component for each bug. This is primarily identified by the associated modular code components in the LLM inference engine. If the relevant faulty code is not within a modular component, they infer its component based on its functionality. For instance, the issue #3821 of vLLM is assigned to the resource manager component. Multi-component bugs are assigned to all relevant components.
Then, inspectors discuss to achieve agreements on the initial codes. In case of disagreements, another author, who has over two years of LLM inference engine deployment experience, acts as an arbitrator and joins the discussion to help reach consensus. After generating the initial codes for issues, inspectors group similar codes into categories. The grouping process is iterative, with inspectors continuously reviewing the categories and issues to refine the taxonomies. Bugs associated with multiple symptoms or root causes are assigned to all relevant categories. Finally, the two inspectors and arbitrator approve all categories in the symptom taxonomy and the root cause taxonomy.